OpenStudy (anonymous):

Find the slope of a line perpendicular to the line that passes through the following points. (3, 9) and (7,15)

OpenStudy (anonymous):

slope of line passing trough these point = 6/4 = 3/2 slope of perpendicular line = -2/3

OpenStudy (anonymous):

ty